cordeliah2007:

A waffle cone has a height of 7 inches and a diameter of 3 inches. What is the volume of ice cream that can be contained within the cone? Use 3.14 for pi. Round your answer to the nearest hundredth.

carmelle:

The formula fo finding the volume of a cone is: \[v = \frac{ 1 }{ 3 }\times h \times \pi \times r^2\] The question already tells us the value of h which is the height, 7 pi is 3.14 (The radius is half of the diameter so basically you do 3 รท 2, which is 1.5) Now all you cave to do is multiply, can you do that? what is v equal to? \[v = \frac{ 1 }{ 3 }\times 7 \times 3.14 \times 1.5^2\]